OpenStudy (anonymous):

Simplify the expression 6n2 - 5n2 + 7n2

OpenStudy (anonymous):

8n^2

OpenStudy (anonymous):

6n2-5n2= n2 n2+7n2=8n2 :D

OpenStudy (anonymous):

thank u seashell :)

OpenStudy (anonymous):

thank u tyler

OpenStudy (anonymous):

youre welcome!:D

OpenStudy (anonymous):

8n2

OpenStudy (anonymous):

6n2 - 5n2 + 7n2 is use 1 variable (n2), then finishing 6-5+7=8, and the end you just add n2